# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 2.3. Good polyhedral decompositions
Figure 2.2. The left hand figure illustrates $\overline{\mathscr{P}}_{1}$ and $\overline{\mathscr{P}}_{2}$. To obtain the final decomposition $\overline{\mathscr{P}}$ the red edges are inserted, as in the right hand figure.
2.4. Maximal degenerations
Given a polarized log Calabi–Yau surface $(Y,D,L)$ with associated Symington polytope $P$ obtained from $\overline{P}$ as in Construction 2.10, we have a natural polyhedral decomposition $\mathscr{P}$ on $P$ obtained by restricting the good polyhedral decomposition $\overline{\mathscr{P}}$ on $\overline{P}$ to $P$. In this section, we describe how to construct a degeneration of $(Y,D,L)$, which is maximal, in the sense that the central fiber has a 00-dimensional stratum. Here, stratum stands for the intersection of irreducible components of the central fiber.
Let $\varphi$ be a convex PL function on $\overline{P}$ defining the regular polyhedral decomposition $\overline{\mathscr{P}}$. We will construct a one-parameter maximal polarized degeneration $(\mathcal{Y}_{\varphi},\mathcal{D}_{\varphi},\mathcal{L}_{\varphi})\rightarrow\mathbb{A}^{1}$, with general fiber deformation to $(Y,D,L)$.
We start with the Mumford toric degeneration $(\overline{\mathcal{Y}}_{\varphi},\overline{\mathcal{D}}_{\varphi},\overline{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}$ of the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ induced by the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$ [39, Example 3.6]. The intersection complex of the central fiber $(\overline{Y}_{\varphi,0},\overline{D}_{\varphi,0},\overline{L}_{\varphi,0})$ is the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$: the irreducible components of the central fiber are the polarized toric surfaces $Y^{f}$ with momentum polytopes the 2-dimensional faces $f$ of $\overline{\mathscr{P}}$ and they are glued along toric divisors $Y^{e}$ corresponding to the edges $e$ of $\overline{\mathscr{P}}$.
Let $h_{\varphi}:\widetilde{\mathcal{Y}}_{\varphi}\rightarrow\overline{\mathcal{Y}}_{\varphi}$ be the blow-up of $\overline{\mathcal{Y}}_{\varphi}$ along the sections $s_{ij}$. We denote by $\widetilde{\mathcal{D}}_{\varphi}$ the strict transform of $\overline{\mathcal{D}}_{\varphi}$, and by ~$\widetilde{\mathcal{L}}_{\varphi}$ the line bundle on $\widetilde{\mathcal{Y}}_{\varphi}$ defined by
~$\widetilde{\mathcal{L}}_{\varphi}=h_{\mathscr{P}}^{\star}\overline{\mathcal{L}}_{\varphi}-\sum_{i=1}^{n}\sum_{j=1}^{n}m_{ij}\mathcal{E}_{ij}\,,$
where $\mathcal{E}_{ij}$ are the exceptional divisors. Comparing with Equation (2.2), it follows that the general fiber of $(\widetilde{\mathcal{Y}}_{\varphi},\widetilde{\mathcal{D}}_{\varphi},\widetilde{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}$ is deformation equivalent to $(\widetilde{Y},\widetilde{D},\widetilde{L})$. While $\widetilde{L}$ is a nef line bundle, the restriction ~$\widetilde{\mathcal{L}}_{\varphi,0}$ of ~$\widetilde{\mathcal{L}}_{\varphi}$ to the central fiber $(\widetilde{\mathcal{Y}}_{\varphi,0},\widetilde{\mathcal{D}}_{\varphi,0})$ is not nef if there exists $i$ and $j$ with $m_{ij}>1$. Indeed, if $m_{ij}>1$, then the exceptional divisor $\mathcal{E}_{ij,0}$ on the central fiber is obtained by blowing-up a point on the divisor $Y^{\overline{P}_{ij}}$, which is contained in the toric surface $Y^{\Delta_{ij,m_{ij}}}_{\overline{\mathscr{P}}}$ with momentum polytope given by the trapezoid $\Delta_{ij,m_{ij}}$ as in Equation (2.5). As $\Delta_{ij,m_{ij}}$ is a trapezoid of size $1$, it follows that $X^{\Delta_{ij,m_{ij}}}_{\overline{\mathscr{P}}}$ is a $\mathbb{P}^{1}$-fibration whose fiber class $F$ satisfies ¯$\overline{\mathcal{L}}_{\varphi,0}\cdot F=1$. After the blow-up, the strict transform of the fiber class passing through the blown-up point is a $(-1)$-curve $C_{ij}$ of class $F-\mathcal{E}_{ij,0}$, and so we have $\widetilde{L}_{\varphi,0}\cdot C_{ij}=1-m_{ij}$, which is negative if $m_{ij}>1$.
@@ 2.4. Maximal degenerations
Figure 2.3. Flop to turn ~$\widetilde{\mathcal{L}}_{\varphi,0}$ into a nef line bundle.
In order to obtain a nef line bundle, we first replace the 3-fold $\widetilde{Y}_{\varphi}$ by its flop along the $(-1,-1)$-curve $C_{ij}$. On the special fiber, this means blowing up the component corresponding to $\Delta_{ij,m_{ij}-1}$ instead of $\Delta_{ij,m_{ij}-1}$. If $m_{ij}>2$, then the line bundle is not nef either, because the intersection with the strict transform of the $\mathbb{P}^{1}$-fiber containing the blown-up point is $2-m_{ij}$, and one can flop this curve – see Figure 2.3. In general, after $m_{ij}-1$ flops, we obtain a family $(\widetilde{\mathcal{Y}}^{\prime}_{\varphi},\widetilde{\mathcal{D}}^{\prime}_{\varphi},\widetilde{\mathcal{L}}^{\prime}_{\varphi})\rightarrow\mathbb{A}^{1}$, whose general fiber is still deformation equivalent to $(\widetilde{Y},\widetilde{D},\widetilde{L})$, but with now the line bundle ~$\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ being relatively big and nef. By [34, Theorem 4.18], a big enough power of ~$\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ is base point free, and so defines a relatively birational contraction
$g:(\widetilde{\mathcal{Y}}^{\prime}_{\varphi},\widetilde{\mathcal{D}}^{\prime}_{\varphi})\rightarrow(\mathcal{Y}_{\varphi},\mathcal{D}_{\varphi})\,,$
and a relatively ample line bundle $\mathcal{L}_{\varphi}$ on $\mathcal{Y}_{\varphi}$ such that ~$\widetilde{\mathcal{L}}^{\prime}_{\varphi}=g^{\star}\mathcal{L}_{\varphi}$. The general fiber of the polarized degeneration
@@
is deformation equivalent to $(Y,D,L)$, and the intersection complex of the central fiber $(\mathcal{Y}_{\varphi,0},\mathcal{D}_{\varphi,0},\mathcal{L}_{\varphi,0})$ is given by the Symington polytope $P$ with its polyhedral decomposition $\mathscr{P}$. In particular, since $\mathscr{P}$ is a decomposition into triangles of size one, it follows that every irreducible component of $\mathcal{Y}_{\varphi,0}$ is isomorphic to $\mathbb{P}^{2}$ and the restriction of $\mathcal{L}_{\varphi,0}$ to each irreducible component of $\mathcal{Y}_{\varphi,0}$ is the line bundle $\mathcal{O}_{\mathbb{P}^{2}}(1)$.
@@ 3. Open Kulikov surfaces and their semistable smoothings
In this section we introduce open Kulikov surfaces and show that $d$-semistable open Kulikov surfaces admit formal semistable smoothings. We also show that these formal smoothings are Mori dream spaces using results of the Minimal Model Program [23] adapted to the formal set up as in [75].
@@ 3.1. Open Kulikov surfaces
Open Kulikov surfaces are normal crossing surfaces whose irreducible components are particular types of surfaces called 00-surfaces, $1$-surfaces, and $2$-surfaces, defined as below.
3.1.1. 00-,$1$-, and $2$-surfaces
@@
An open log Calabi–Yau surface $(Y,D)$ is a smooth complex surface $Y$, together with a non-empty reduced normal crossing divisor $D\subset Y$ such that there exists a smooth log Calabi–Yau surface $(\widetilde{Y},\widetilde{D})$ and a non-empty connected curve $Z$ in $\widetilde{Y}$ given by a union of irreducible components of $\tilde{D}$, such that $Y=\widetilde{Y}\setminus Z$ and $D=\tilde{D}\setminus Z$.
Let $(Y,D)$ be an open log Calabi–Yau surface as in Definition 3.1. Then, for a compactification $(\widetilde{Y},\widetilde{D})$ of $(Y,D)$ as in Definition 3.1, the divisor $\widetilde{D}$ has at least two irreducible components, and so is a cycle of projective lines. It follows that the divisor $D$ is a chain $D_{0}+\cdots+D_{n}$, where $D_{0}\simeq D_{n}\simeq\mathbb{A}^{1}$, and $D_{i}\simeq\mathbb{P}^{1}$ for $11\leq i\leq n-$ 1.
@@
A 00-surface is an open log Calabi–Yau surface $(Y,D)$ obtained by interior and corner blow-ups of an affine toric surface $(\overline{Y},\overline{D})$, where the interior blow-ups are all done above the torus fixed point of $\overline{Y}$.
If $(Y,D)$ is a 00-surface, then one can uniquely recover $(\overline{Y},\overline{D})$ as in Definition 3.2 as the affinization of $(Y,D)$, that is,
@@
and the affinization morphism $(Y,D)\rightarrow(\overline{Y},\overline{D})$ is projective, surjective, and birational. In particular, $Y$ contains only finitely many compact irreducible curves, which are either compact irreducible components of $D$, or irreducible components of generalized exceptional curves. Moreover, $\overline{Y}$ is a cyclic quotient singularity, $Y\rightarrow\overline{Y}$ is a resolution of singularities, and the union of compact curves in $Y$ is the corresponding exceptional divisor. In particular, we have $C^{2}<0$ for every compact curve in $Y$.
@@
The fan of the affine toric surface $(\overline{Y},\overline{D})$ is a strictly convex rational cone $\sigma$ in $\mathbb{R}^{2}$. A toric 00-surface is then a toric surface whose fan is a refinement of $\sigma$. Finally, any 00-surface $(Y,D)$ is obtained from a toric 00-surface $(Y^{\prime},D^{\prime})$ by interior blow-ups. The tropicalization of $(Y,D)$ is obtained by introducing focus-focus singularities moving along the interior rays of the fan of $(Y^{\prime},D^{\prime})$, and pushing them to the origin, as in [41].
@@
A $1$-surface is an open log Calabi–Yau surface $(Y,D)$ such that there exists a projective surjective morphism $Y\rightarrow\mathbb{A}^{1}$, contracting the compact irreducible components of $D$ to $0\in\mathbb{A}^{1}$.
If $(Y,D)$ is a $1$-surface, then one can recover $(\mathbb{A}^{1},0)$ as the affinization of $(Y,D)$, that is,
$(\mathbb{A}^{1},0):=(\mathrm{Spec}\,H^{0}(Y,\mathcal{O}_{Y}),\mathrm{Spec}\,H^{0}(D,\mathcal{O}_{D}))\,.$
Moreover, the natural projection $(Y,D)\rightarrow(\mathbb{A}^{1},0)$ is the affinization morphism, and is projective, surjective, and generically a $\mathbb{P}^{1}$-fibration. Moreover, every compact irreducible curve of $Y$ is contained in a fiber of $(Y,D)\rightarrow(\mathbb{A}^{1},0)$, and so $Y$ contains a $1$-dimensional family of compact curves, but no $2$-dimensional family of compact curves. In particular, we have $C^{2}\leq 0$ for every compact curve in $Y$, and we have $C^{2}=0$ if and only if $C$ is a fiber of the ruling $Y\rightarrow\mathbb{A}^{1}$.
@@
Note that a $2$-surface is projective, and so is covered by families of compact curves of dimension $\geq 2$.
@@ 3.1.2. Open Kulikov surfaces
    Let $\{X^{i}\}_{i\in I}$ be the set of irreducible components of $\mathcal{X}_{0}$, and denote by $\{v_{i}\}_{i\in I}$ the corresponding vertices in $B$. For every $i\in I$, denote the union of double curves in $X^{i}$ by $\partial X^{i}=\cup_{j\in I\setminus\{i\}}X^{i}\cap X^{j}$. Denote by $\tilde{X}^{i}$ the normalization of $X^{i}$ and by $\partial\tilde{X}^{i}$ the pre-image of $\partial X^{i}$ in $\tilde{X}^{i}$. Then, for each interior vertex $v_{i}\in\mathrm{Int}(B)$, the pair $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is a 2-surface, and for each boundary vertex $v_{i}\in\partial B$, the pair $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is either a 00-surface or a $1$-surface.
@@
    The surface $\mathcal{X}_{0}$ is combinatorially $d$-semistable, that is, for every irreducible component $D_{ij}=X^{i}\cap X^{j}$, we have
    $(D_{ij}|_{X^{i}})^{2}+(D_{ij}|_{X_{j}})^{2}=-2+2g(D_{ij})\,,$
    where $g(D_{ij})$ is the arithmetic genus of $D_{ij}$.
@@
The dual intersection complex $B$ of an open Kulikov surface $\mathcal{X}_{0}$ is a simplicial complex, containing a vertex for every irreducible component, a 1-simplex for each irreducible component of the double locus, and a 2-simplex for each triple point. By Definition 3.7 i), the topological realization $P$ of $B$ is homeomorphic to a disk, and we denote by $\mathscr{P}$ the triangulation into triangles of integral size $1$ induced by the simplicial structure. Moreover, it follows from Definition 3.7 iii) that the natural integral affine structure on the interior of the triangles of $\mathscr{P}$ can be naturally extended to the complement of the vertices of $\mathscr{P}$. This integral affine structure extends over a vertex $v_{i}$ if the corresponding irreducible component $(X^{i},\partial X^{i})$ is a toric pair, but not in general. We refer to $(P,\mathscr{P})$ as the tropicalization of $\mathcal{X}_{0}$, where $P$ is viewed as an integral affine disk with singularities.
@@ 3.1.3. Properties of open Kulikov surfaces
the corresponding resolution of $\mathcal{O}_{\mathcal{X}_{\mathscr{P},0}}^{\star}$, and the fact that the tropicalization $P$ is topologically a disk and so $H^{0}(P,\mathbb{Z})=0$, and $H^{i}(P,\mathbb{Z})=0$ for $i>0$.

### v3

@@ 2.4. Maximal degenerations
Figure 4. Flop to turn $\widetilde{\mathcal{L}}_{\varphi,0}$ into a nef line bundle.
@@
$g:(\widetilde{\mathcal{Y}}^{\prime}_{\varphi},\widetilde{\mathcal{D}}^{\prime}_{\varphi})\longrightarrow(\mathcal{Y}_{\varphi},\mathcal{D}_{\varphi})\,$,
and a relatively ample line bundle $\mathcal{L}_{\varphi}$ on $\mathcal{Y}_{\varphi}$ such that $\widetilde{\mathcal{L}}^{\prime}_{\varphi}=g^{\star}\mathcal{L}_{\varphi}$. By construction, the line bundle $\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ is trivial in restriction to every double curve of the central fiber $\widetilde{\mathcal{Y}}^{\prime}_{\varphi,0}$ corresponding to an edge of $\overline{\mathscr{P}}$ contained in the interior of triangles $\Delta_{ij}$. Moreover, $\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ agrees with $\overline{\mathcal{L}}_{\varphi}$, and so is ample, on every irreducible component of $\widetilde{\mathcal{Y}}^{\prime}_{\varphi,0}$ corresponding to a 2-dimensional face of $\overline{\mathscr{P}}$ not contained in the triangles $\Delta_{ij}$. Hence, the irreducible components of $\widetilde{\mathcal{Y}}^{\prime}_{\varphi,0}$ contracted by $g$ are exactly the ones corresponding to the 2-dimensional faces of $\overline{\mathscr{P}}$ contained in the triangles $\Delta_{ij}$. Thus, the general fiber of the polarized degeneration
@@
is deformation equivalent to $(Y,D,L)$, and the intersection complex of the central fiber $(\mathcal{Y}_{\varphi,0},\mathcal{D}_{\varphi,0},\mathcal{L}_{\varphi,0})$ is given by the Symington polytope $P$ with its polyhedral decomposition $\mathscr{P}$. In particular, since $\mathscr{P}$ is a decomposition into triangles of size one, the normalization of every irreducible component of $\mathcal{Y}_{\varphi,0}$ is isomorphic to $\mathbb{P}^{2}$, and the restriction of $\mathcal{L}_{\varphi,0}$ to the normalization of each irreducible component of $\mathcal{Y}_{\varphi,0}$ is the ample line bundle $\mathcal{O}_{\mathbb{P}^{2}}(1)$.
@@ 3. Open Kulikov surfaces and their semistable smoothings
In this section, we introduce open Kulikov surfaces and establish that $d$-semistable open Kulikov surfaces admit formal semistable smoothings. We also show that these formal smoothings are Mori dream spaces using results of the Minimal Model Program [24] adapted to the formal setup as in [80].
@@ 3.1. Open Kulikov surfaces
Open Kulikov surfaces are normal crossing surfaces whose irreducible components are particular types of surfaces called 0-surfaces, 11-surfaces, and $2$-surfaces, defined as below.
3.1.1. 0-,11-, and $2$-surfaces
@@
[…]
@@
A 0-surface is an open log Calabi–Yau surface $(Y,D)$ such that, denoting
@@
the affinization of $(Y,D)$, and $\pi:(Y,D)\rightarrow(\overline{Y},\overline{D})$ the affinization morphism, the following holds:
  • i)
    $(\overline{Y},\overline{D})$ is an affine toric surface with a torus fixed point $x_{0}$.
  • ii)
    The morphism $\pi$ is birational, projective, and surjective.
  • iii)
    Every $\pi$-exceptional curve is contracted by $\pi$ to $x_{0}$.
If $(Y,D)$ is a 0-surface, then $Y$ contains only finitely many compact irreducible curves. Moreover, $\overline{Y}$ is a cyclic quotient singularity, $Y\rightarrow\overline{Y}$ is a resolution of singularities, and the union of compact curves in $Y$ is the corresponding exceptional divisor. In particular, we have $C^{2}<0$ for every compact curve in $Y$.
@@
The fan of the affine toric surface $(\overline{Y},\overline{D})$ is a strictly convex rational cone $\sigma$ in $\mathbb{R}^{2}$. A toric 0-surface is then a toric surface whose fan is a refinement of $\sigma$. Finally, any 0-surface $(Y,D)$ is obtained from a toric 0-surface $(Y^{\prime},D^{\prime})$ by interior blow-ups. The tropicalization of $(Y,D)$ is obtained by introducing focus-focus singularities moving along the interior rays of the fan of $(Y^{\prime},D^{\prime})$, and pushing them to the origin, as in [44].
@@
A $1$-surface is an open log Calabi–Yau surface $(Y,D)$ with affinization $\mathrm{Spec}\,H^{0}(Y,\mathcal{O}_{Y})\simeq\mathbb{A}^{1}$, such that, denoting by $\pi:Y\rightarrow\mathbb{A}^{1}$ the affinization morphism and $0\in\mathbb{A}^{1}$ the image by $\pi$ of the compact components of $D$, the following holds:
  • i)
    The morphism $\pi$ is a projective surjective morphism.
  • ii)
    $\pi^{-1}(\mathbb{A}^{1}\setminus\{0\})\simeq\mathbb{P}^{1}\times(\mathbb{A}^{1}\setminus\{0\})$.
  • iii)
    $\pi|_{\pi^{-1}(\mathbb{A}^{1}\setminus\{0\})}:\mathbb{P}^{1}\times(\mathbb{A}^{1}\setminus\{0\})\rightarrow(\mathbb{A}^{1}\setminus\{0\})$ is the projection on the second factor.
If $(Y,D)$ is a $1$-surface, then every compact curve of $Y$ is contained in a fiber of $Y\rightarrow\mathbb{A}^{1}$. This implies that $Y$ contains a $1$-dimensional family of compact curves, but does not contain a $2$-dimensional family of compact curves. In particular, we have $C^{2}\leq 0$ for every compact curve $C$ in $Y$, and we have $C^{2}=0$ if and only if $C$ is a fiber of $\pi:Y\rightarrow\mathbb{A}^{1}$.
@@
A $2$-surface is projective, and so is covered by families of compact curves of dimension $\geq 2$.
@@ 3.1.2. Open Kulikov surfaces
    Let $\{X^{i}\}_{i\in I}$ be the set of irreducible components of $\mathcal{X}_{0}$, and denote by $\{v_{i}\}_{i\in I}$ the corresponding vertices in $B$. For every $i\in I$, denote the union of double curves in $X^{i}$ by $\partial X^{i}$. Denote by $\tilde{X}^{i}$ the normalization of $X^{i}$ and by $\partial\tilde{X}^{i}$ the pre-image of $\partial X^{i}$ in $\tilde{X}^{i}$. Then, for each interior vertex $v_{i}\in\mathrm{Int}(B)$, the pair $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is a 2-surface, and for each boundary vertex $v_{i}\in\partial B$, the pair $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is either a 0-surface or a $1$-surface. Moreover, there exists at least one boundary vertex $v_{i}\in\partial B$ such that the pair $(\tilde{X}^{i},\partial\tilde{X}^{i})$ is a 0-surface.
@@
    The surface $\mathcal{X}_{0}$ is combinatorially $d$-semistable, that is, for every proper irreducible double curve $C$ of $\mathcal{X}_{0}$, denoting by $C_{i}\subset\tilde{X}^{i}$ and $C_{j}\subset\tilde{X}_{j}$ the two connected components of the pre-image of $C$ in the normalization of $\mathcal{X}_{0}$, we have
    $(C_{i}|_{\tilde{X}^{i}})^{2}+(C_{j}|_{\tilde{X}_{j}})^{2}=-2+2g(C)\,$,
    where $g(C)$ is the arithmetic genus of $C$.
@@
The dual intersection complex $B$ of an open Kulikov surface $\mathcal{X}_{0}$ is a simplicial complex, containing a vertex for every irreducible component, a 1-simplex for each irreducible component of the double locus, and a 2-simplex for each triple point. By Definition 3.7 i), the topological realization $P$ of $B$ is homeomorphic to a disk, and we denote by $\mathscr{P}$ the triangulation into triangles of integral size $1$ induced by the simplicial structure. Furthermore, Definition 3.7 iii) implies that the natural integral affine structure on the interior of the triangles of $\mathscr{P}$ can be naturally extended to the complement of the vertices of $\mathscr{P}$. This integral affine structure extends over a vertex $v_{i}$ if the corresponding irreducible component $(X^{i},\partial X^{i})$ is a toric pair, but not in general. We refer to $(P,\mathscr{P})$ as the tropicalization of $\mathcal{X}_{0}$, where $P$ is viewed as an integral affine disk with singularities.
@@ 3.1.3. Properties of open Kulikov surfaces
the corresponding resolution of $\mathcal{O}_{\mathcal{X}_{\mathscr{P},0}}^{\star}$, and the fact that the tropicalization $P$ is topologically a disk and so $H^{0}(P,\mathbb{Z})=\mathbb{Z}$, and $H^{i}(P,\mathbb{Z})=0$ for $i>0$. ∎
Lemma 3.11.
Let $\mathcal{X}_{0}$ be an open Kulikov surface with tropicalization $(P,\mathscr{P})$. Then, we have $\mathrm{rk}\mathrm{Pic}(\mathcal{X}_{0})\geq Q+|P_{\mathbb{Z}}|-3$, where $Q$ is the charge of $(Y,D)$ as in [36, Definition 1.1], that is, the number of interior blow-ups in a toric model of $(Y,D)$, and $|P_{\mathbb{Z}}|$ is the number of integral points in $P$.
Proof.
Denote by $V=|P_{\mathbb{Z}}|$, $E$, $E_{\mathrm{int}}$, $F$ the number of vertices, edges, interior edges, and faces of $\mathscr{P}$ respectively. For every verter $v$, we have $\mathrm{rk}\mathrm{Pic}(X^{v})=Q^{v}+E^{v}-2$, where $Q^{v}$ is the charge of $X^{v}$, and $E^{v}$ is the number of edges adjacent to $v$. Thus, by Lemma 3.10, we obtain
$\mathrm{rk}\,\mathrm{Pic}(\mathcal{X}_{0})\geq\left(\sum_{v\in\mathscr{P}^{[0]}}Q_{v}\right)+\left(\sum_{v\in\mathscr{P}^{[0]}}E^{v}\right)-2V-E_{int}=Q+2E-2V-E_{int}\,.$
As $\mathscr{P}$ is a triangulation of a disk, we have VE+F=1V-E+F=1 and $3F=E+E_{int}$, and so
$2E-2V-E_{int}=2E-2V-(3F-E)=2E-2V-3(1+E-V)+E=V-3\,$,
Remark 3.12.
We will see in Corollary 8.14 that we actually have $\mathrm{rk}\mathrm{Pic}(\mathcal{X}_{0})=Q+|P_{\mathbb{Z}}|-3$.